# How Safe Will I Be Given What I Saw? Calibrated Prediction of Safety Chances for Image-Controlled Autonomy

### v2

@@ 5 Experimental Evaluation
Benchmark Systems.
The racing car and cart pole from the OpenAI Gym Brockman et al. (2016) and the realistic donkey car are selected as our study cases. These environments are more suitable than pre-collected robotic datasets, such as the Waymo Open Dataset, for two reasons. First, to study safety prediction, we require a large dataset of safety violations, which is rarely found in real-world data. Second, the simulated dynamical systems are convenient for collecting an unbounded amount of data with direct access to the images and the ground truth for experimental evaluation (e.g., the car’s position).
Performance Metrics.
We use the F1 score as our main metric for evaluating the accuracy of safety label predictors, as it balances the precision and recall: $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}=2\times\frac{TP}{{TP+\frac{FP+FN}{2}}}$. False positives are also a major concern in safety prediction (actually unsafe situations predicted as safe), so we also evaluate the False Positive Rate (FPR): $\text{FPR}=\frac{FP}{FP+TN}$. To evaluate our chance predictors, we compute the Expected Calibration Error (ECE, intuitively a weighted average difference between the predicted and true probabilities) and the Brier Score (mean squared error between predicted probabilities and actual binary outcomes) Guo et al. (2017); Minderer et al. (2021); Glenn et al. (1950).
Table 3: F1 Score and FPR.
Model Racing Car Cart Pole Donkey Car
k=60 k=120 k=180 k=10 k=20 k=30 k=30 k=60 k=90
F1 score
mono. CNN 88.04 79.51 73.83 99.65 94.45 95.21 88.99 83.73 77.18
mono. LSTM 87.89 80.48 74.95 98.29 93.97 85.72 84.69 79.92 69.89
mono. BiLSTM 86.88 80.24 75.09 98.59 91.35 86.76 83.69 78.72 72.27
mono. Transformer 87.34 80.89 74.55 98.38 94.45 87.21 88.99 77.83 73.25
comp. LSTM 85.99 73.67 73.60 99.65 99.11 87.50 85.87 82.66 76.19
comp. BiLSTM 90.17 78.93 61.54 99.65 98.37 87.02 87.49 80.55 76.73
comp. Transformer 83.83 76.62 70.18 99.47 98.93 95.21 84.46 76.03 74.59
comp. VQ-Transformer 94.55 89.35 79.03 98.65 93.83 88.21 88.18 83.73 77.18
FPR
mono. CNN 27.25 39.48 54.11 8.41 28.43 34.39 26.35 37.16 42.83
mono. LSTM 26.57 35.30 44.56 5.31 24.12 36.44 24.89 35.16 38.66
mono. BiLSTM 26.69 36.35 45.34 10.36 22.27 35.93 27.47 33.63 47.54
mono. Transformer 25.48 36.81 42.80 13.86 34.08 37.29 25.76 31.29 40.33
comp. LSTM 16.75 23.98 35.41 10.38 19.68 39.45 32.56 30.46 36.71
comp. BiLSTM 28.57 30.71 34.29 3.09 16.92 46.33 28.87 27.23 40.37
comp. Transformer 17.13 26.56 35.71 14.51 30.91 31.34 41.10 30.36 47.90
comp. VQ-Transformer 18.18 27.01 35.72 3.56 24.94 28.28 32.12 34.73 41.92
@@ 5.2 Comparative Results
Result 1: Composite predictors outperform monolithic ones.
Monolithic predictors use their full learning capacity to predict safety but do not learn the underlying dynamics. Therefore, our initial hypothesis was that they would do well on short horizons — but lose to composite predictors on longer horizons. The results support this hypothesis, as illustrated in Figure 4: the performance of monolithic predictors degrades faster as the horizon grows. Moreover, composite predictors offered a better FPR as shown in Table 3 for the donkey cars, which may be desirable in safety-critical systems. We contend that the inferior performance of monolithic predictors arises from the inherent difficulty of learning coherent, long‑term latent dynamics — a challenge that remains largely unresolved. By contrast, composite predictors achieve superior results by explicitly decomposing the safety‐prediction problem into three stages: (1) learning a suitable latent representation; (2) forecasting future states, and (3) evaluating those states’ safety. This architectural separation imposes a clear inductive bias: the model is first guided to generate plausible trajectories and then to judge their safety, rather than expecting a single network to discover both the underlying dynamics and the safety criterion simultaneously. As a result, each subtask becomes substantially more tractable, leading to higher overall performance.
Result 2: Attention mechanisms significantly improve composite predictor performance.
@@
Result 3: Latent evaluators outperform image-based ones.
Latent evaluators outperform image-based ones both in F1 and FPR, as shown in the Tables 1 and 2. This is primarily because the VAE’s reconstruction objective encourages the latent space to capture fine-grained image details, while the prior imposes additional structure that facilitates classification. Moreover, KL regularization promotes smoothness of the latent manifold and can implicitly disentangle task-relevant features. Additionally, the latent representation filters out pixel-level noise and, by mapping sequential observations into a compact and structured manifold, preserves the temporal relationships between states. This smoothness in latent trajectories makes it easier for the evaluator to capture long-term safety-relevant dynamics and avoid being misled by transient visual variations.
Result 4: UDA evaluator improves robustness under distribution shift.
We evaluate the effectiveness of unsupervised domain adaptation in enhancing the evaluator’s robustness to distribution shift. In addition to the normal images, we test on two special test subsets: OOD, which consists of 500 images sampled from predictions with manual labels that exhibit strong distribution shift, and Mix, which randomly selects 500 predictions containing both in-distribution and OOD examples, with proportions varying by case study.. As shown in Tables 1 and 2, across all case studies, UDA-based image evaluators consistently outperform both standard image and latent evaluators on the OOD and Mix subsets, without losing almost any performance on the normal data. These results demonstrate that UDA is a promising technique to generalize safety evaluators to challenging out-of-distribution scenarios.
Result 5: Calibration improves safety‐chance predictor reliability.
@@
Result 6: Conformal calibration coverage is reliable.
@@ 6 Discussion and Conclusion
Limitations: our predictor family’s scope is limited to systems where safety can be inferred from raw sensor data. This assumption holds in domains like autonomous driving, where an imminent collision, lane departure, or off-track event can be visually detected. In contrast, there exist systems where safety depends on unobservable internal states or variables that are not visually apparent. For example, for an aircraft whose structural integrity depends on internal component stresses, raw visual or range data would be insufficient to assess safety. Besides, obtaining negative safety samples can have a high cost in real-world systems; however, this issue can be mitigated via transfer learning from simulation and generative models.
Some approaches have performed poorly, like monolithic flexible-horizon predictors, which output the time to the first expected safety violation. Their prediction space appeared too complex and insufficiently regularized, leading to poor performance. Also, existing test-time adaptation methods operate at the image level, which is potentially inefficient. Exploring adaptation in latent representation could further improve the performance of the latent evaluator under distribution shift. Thus, learning strong safety-informed representations for autonomy remains an open problem. With the rise of large language models, meaning representations in multimodal autoregressive models Liu et al. (2023) may provide more possibilities to explore the surrogate dynamics of autonomy systems.
Future work includes adding physical constraints to latent states similar to Neural ODEs Wen et al. (2022), using a generative model to obtain surrogate unsafe samples from a more realistic environment, and applying vision-language and vision-language-action models to automatically label the safety of generative predictions.
@@
{acks}
{dci}
The authors declared no potential conflicts of interest with respect to the research, authorship, or publication of this article.

### v3

@@ 5 Experimental Evaluation
Benchmark Systems
The racing car and cart pole from the OpenAI Gym and the realistic donkey car are selected as our study cases [brockman_openai_2016]. These environments are more suitable than pre-collected robotic datasets, such as the Waymo Open Dataset, for two reasons. First, to study safety prediction, we require a large dataset of safety violations, which is rarely found in real-world data. Second, the simulated dynamical systems are convenient for collecting an unbounded amount of data with direct access to the images and the ground truth for experimental evaluation (e.g., the car’s position).
Performance Metrics
We use the F1 score as our main metric for evaluating the accuracy of safety label predictors, as it balances the precision and recall:
$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}=2\times\frac{TP}{TP+\frac{FP+FN}{2}}.$
False positives are also a major concern in safety prediction (actually unsafe situations predicted as safe), so we also evaluate the False Positive Rate (FPR):
$\text{FPR}=\frac{FP}{FP+TN}.$
To evaluate our chance predictors, we compute the Expected Calibration Error (ECE, intuitively a weighted average difference between the predicted and true probabilities) and the Brier Score (mean squared error between predicted probabilities and actual binary outcomes) [guo_calibration_2017, minderer_revisiting_2021, glenn1950verification].
@@ 5.2 Comparative Results
Table 1: F1 Score and FPR.
Model Racing Car Cart Pole Donkey Car k=60 k=120 k=180 k=10 k=20 k=30 k=30 k=60 k=90 F1 score mono. CNN 88.04 79.51 73.83 99.65 94.45 95.21 88.99 83.73 77.18 mono. LSTM 87.89 80.48 74.95 98.29 93.97 85.72 84.69 79.92 69.89 mono. BiLSTM 86.88 80.24 75.09 98.59 91.35 86.76 83.69 78.72 72.27 mono. Transformer 87.34 80.89 74.55 98.38 94.45 87.21 88.99 77.83 73.25 comp. LSTM 85.99 73.67 73.60 99.65 99.11 87.50 85.87 82.66 76.19 comp. BiLSTM 90.17 78.93 61.54 99.65 98.37 87.02 87.49 80.55 76.73 comp. Transformer 83.83 76.62 70.18 99.47 98.93 95.21 84.46 76.03 74.59 comp. VQ-Transformer 94.55 89.35 79.03 98.65 93.83 88.21 88.18 83.73 77.18 FPR mono. CNN 27.25 39.48 54.11 8.41 28.43 34.39 26.35 37.16 42.83 mono. LSTM 26.57 35.30 44.56 5.31 24.12 36.44 24.89 35.16 38.66 mono. BiLSTM 26.69 36.35 45.34 10.36 22.27 35.93 27.47 33.63 47.54 mono. Transformer 25.48 36.81 42.80 13.86 34.08 37.29 25.76 31.29 40.33 comp. LSTM 16.75 23.98 35.41 10.38 19.68 39.45 32.56 30.46 36.71 comp. BiLSTM 28.57 30.71 34.29 3.09 16.92 46.33 28.87 27.23 40.37 comp. Transformer 17.13 26.56 35.71 14.51 30.91 31.34 41.10 30.36 47.90 comp. VQ-Transformer 18.18 27.01 35.72 3.56 24.94 28.28 32.12 34.73 41.92
@@
Composite predictors outperform monolithic ones
Monolithic predictors use their full learning capacity to predict safety but do not learn the underlying dynamics. Therefore, our initial hypothesis was that they would do well on short horizons — but lose to composite predictors on longer horizons. The results support this hypothesis, as illustrated in Figure 5: the performance of monolithic predictors degrades faster as the horizon grows. Moreover, composite predictors offered a better FPR as shown in Table 1 for the donkey cars, which may be desirable in safety-critical systems. We contend that the inferior performance of monolithic predictors arises from the inherent difficulty of learning coherent, long-term latent dynamics — a challenge that remains largely unresolved. By contrast, composite predictors achieve superior results by explicitly decomposing the safety‐prediction problem into three stages: (1) learning a suitable latent representation; (2) forecasting future states, and (3) evaluating those states’ safety. This architectural separation imposes a clear inductive bias: the model is first guided to generate plausible trajectories and then to judge their safety, rather than expecting a single network to discover both the underlying dynamics and the safety criterion simultaneously. As a result, each subtask becomes substantially more tractable, leading to higher overall performance.
Table 3: F1 scores of evaluators on different datasets (in %).
Case Study Evaluator Type Normal OOD Mix
Racing Car Image Evaluator 99.56 63.94 78.23
Latent Evaluator 98.79 65.73 80.27
UDA Image Evaluator 97.13 81.65 91.34
Cart Pole Image Evaluator 99.66 67.21 91.05
Latent Evaluator 99.67 69.14 89.33
UDA Image Evaluator 99.06 76.52 96.62
Donkey Car Image Evaluator 96.20 93.03 94.43
Latent Evaluator 97.19 92.21 94.70
UDA Image Evaluator 96.05 94.14 95.97
Table 4: FPR of evaluators on different datasets (in %).
Case Study Evaluator Type Normal OOD Mix
Racing Car Image Evaluator 0.42 52.78 23.53
Latent Evaluator 2.15 30.82 19.62
UDA Image Evaluator 0.62 7.07 4.37
Cart Pole Image Evaluator 0.53 21.51 9.88
Latent Evaluator 0.61 18.34 7.69
UDA Image Evaluator 1.34 16.72 10.03
Donkey Car Image Evaluator 2.69 18.25 6.05
Latent Evaluator 3.05 11.97 5.32
UDA Image Evaluator 3.21 9.15 5.42
Attention significantly improves composite predictor performance
@@
Latent evaluators outperform image-based ones
Latent evaluators outperform image-based ones both in F1 and FPR, as shown in the Tables 3 and 4. This is primarily because the VAE’s reconstruction objective encourages the latent space to capture fine-grained image details, while the prior imposes additional structure that facilitates classification. Moreover, KL regularization promotes smoothness of the latent manifold and can implicitly disentangle task-relevant features. Additionally, the latent representation filters out pixel-level noise and, by mapping sequential observations into a compact and structured manifold, preserves the temporal relationships between states. This smoothness in latent trajectories makes it easier for the evaluator to capture long-term safety-relevant dynamics and avoid being misled by transient visual variations.
UDA evaluator improves robustness under distribution shift
We evaluate the effectiveness of unsupervised domain adaptation in enhancing the evaluator’s robustness to distribution shift. In addition to the normal images, we test on two special test subsets: OOD, which consists of 500 images sampled from predictions with manual labels that exhibit strong distribution shift, and Mix, which randomly selects 500 predictions containing both in-distribution and OOD examples, with proportions varying by case study. As shown in Tables 3 and 4, across all case studies, UDA-based image evaluators consistently outperform both standard image and latent evaluators on the OOD and Mix subsets, without losing almost any performance on the normal data. These results demonstrate that UDA is a promising technique to generalize safety evaluators to challenging out-of-distribution scenarios.
Calibration improves safety‐chance predictor reliability.
@@
Conformal calibration coverage is reliable
@@ 6 Discussion and Conclusion
Limitations
Our predictor family’s scope is limited to systems where safety can be inferred from raw sensor data. This assumption holds in domains like autonomous driving, where an imminent collision, lane departure, or off-track event can be visually detected. Relatedly, the safety specifications considered in this work are restricted to simple state-based binary predicates ($\varphi:\mathcal{X}\rightarrow\{0,1\}$) and do not capture temporal requirements. In contrast, there exist systems where safety depends on unobservable internal states or variables that are not visually apparent. For example, for an aircraft whose structural integrity depends on internal component stresses, raw visual or range data would be insufficient to assess safety.
Obtaining labeled safety data is associated with notable costs, especially for negative safety samples in physical systems. Our approach is limited to systems where it is possible to generate a large volume of labeled data to effectively train the predictors. For instance, our donkey car case study required labeling over 10,000 images, which took approximately 5-10 hours of human effort, including iterative refinement. While this labeling burden is non-trivial, it is feasible for many robotics applications where safety-critical scenarios can be systematically collected. Furthermore, this burden can be substantially alleviated through automated and semi-automated labeling techniques. Recent frameworks for active learning, weak supervision, and programmatic labeling have demonstrated significant reductions in manual annotation effort for similar vision-based tasks [settles2009active, ratner2017snorkel, ratner2019snorkel].
Some prediction approaches have performed poorly, like monolithic flexible-horizon predictors, which output the time to the first expected safety violation. Their prediction space appeared too complex and insufficiently regularized, leading to poor learning and performance. Also, existing test-time adaptation methods operate at the image level, which is potentially inefficient. Exploring adaptation in latent representation could further improve the performance of the latent evaluator under distribution shift. Thus, learning strong safety-informed representations for autonomy remains an open problem. With the rise of large language models, meaning representations in multimodal autoregressive models may provide more possibilities to explore the surrogate dynamics of autonomy systems [liu2023meaning].
Future work
Interesting future directions include adding physical constraints to latent states similar to Neural ODEs, using a generative model to obtain surrogate unsafe samples from a more realistic environment, and applying vision-language and vision-language-action models to automatically label the safety of generative predictions [wen_social_2022]. Another opportunity is to mitigate the labeling burden via transfer learning from simulation and generative models. Furthermore, our framework can be extended to support safety specifications in linear temporal logic (LTL), enabling safety predictions that reason over sequences of events and long-horizon behavioral constraints [hahn2021teaching].
Declarations
Acknowledgments